How much will a firm receive in net funding from an underwriting of 100,000 shares priced to the public at $30 if an 8% underwriting spread has been paid to the underwriter? Additionally, the firm pays $600,000 in legal fees
Answer #1

The total amount receivable from issue of shares is (100,000*$30) = $3,000,000.

The expenses for the same is 8% underwriting fee and $600,000 legal fee.

Thus, the net amount receivable is:

Net\, amount=\left [3,000,000\times \left ( 1-0.08 \right ) \right ]-600,000

Net\, amount=\left [3,000,000\times 0.92 \right ]-600,000

Net\, amount=2,760,000-600,000

Net\, amount=2,160,000

The net amount receivable is $2,160,000.